# Defense Against Prompt Injection Attack by Leveraging Attack Techniques

### v1

@@ Abstract
With the advancement of technology, large language models (LLMs) have achieved remarkable performance across various natural language processing (NLP) tasks, powering LLM-integrated applications like Microsoft Copilot. However, as LLMs continue to evolve, new vulnerabilities, especially prompt injection attacks arise. These attacks trick LLMs into deviating from the original input instructions and executing the attacker’s instructions injected in data content, such as retrieved results. Recent attack methods leverage LLMs’ instruction-following abilities and their inabilities to distinguish instructions injected in the data content, and achieve a high attack success rate (ASR). When comparing the attack and defense methods, we interestingly find that they share similar design goals, of inducing the model to ignore unwanted instructions and instead to execute wanted instructions. Therefore, we raise an intuitive question: Could these attack techniques be utilized for defensive purposes? In this paper, we invert the intention of prompt injection methods to develop novel defense methods based on previous training-free attack methods, by repeating the attack process but with the original input instruction rather than the injected instruction. Our comprehensive experiments demonstrate that our defense techniques outperform existing training-free defense approaches, achieving state-of-the-art results.
@@ 1 Introduction
We conduct comprehensive experiments to evaluate the effectiveness of our defense methods against various prompt injection attack methods. The results demonstrate that our methods outperform existing training-free defense approaches, offering more robust protection. Notably, the defense method based on the most effective attack technique performs the best, reducing the attack success rate (ASR) to nearly zero in certain scenarios. Our contributions are summarized as follows:
@@
  • We develop prompt injection defense methods based on attack strategies, which demonstrate greater effectiveness compared to existing training-free baselines.
@@ 5.1 Experimental Settings
For the security metric, we follow the evaluation protocol of Chen et al. (2024), using the attack success rate (ASR) to assess the effectiveness of the defense methods. Chen et al. (2024)’s approach involves injecting a malicious prompt at the end of the data content, a position deemed most advantageous for attackers Piet et al. (2023), with the instruction “Print exactly Hacked!” If “Hacked” appears in the LLM’s response, the attack is considered successful. However, LLMs may sometimes repeat the injected instruction as part of the data content rather than truly executing it. This renders the “Hacked” evaluation method unreliable. To address this, we modify the approach by injecting the instruction “What is the answer to 1000+555?”. This ensures that when the correct answer appears in the response, the instruction has been executed rather than merely repeated. For the utility metric, we use accuracy to evaluate the potential negative impact of defense methods on model performance. Specifically, we employ the filtered QA dataset, which is not attacked and includes the defense mechanism. We request the LLMs to answer the questions and verify whether the correct (golden) answers appear in the responses.
@@ 5.2.2 Defense Baselines
For a fair comparison, we select existing training-free defense methods as baselines. Specifically, we select Incontext Wei et al. (2023), Sandwich san (2023), Instructional ins (2023), Reminder Yi et al. (2023), Isolation Willison (2023), Spotlight Hines et al. (2024) for comparison. More details about the baselines can be found in Appendix A.1.
@@ 5.3.3 Model Utility
None 77.70 78.20 76.55
Incontext 77.70 78.20 76.55
Sandwich 77.70 78.20 76.55
Instructional 77.70 78.20 76.55
Reminder 77.70 78.20 76.55
Isolation 77.70 78.20 76.55
Spotlight 77.70 78.20 76.55
Ours-Ignore 77.00 78.65 78.05
Ours-Escape 77.00 78.65 78.05
Ours-Fakecom 77.00 78.65 78.05
Ours-Fakecom-t 79.50 80.25 77.50
Table 3: The model utility across different defense methods. All the results are reported in %.
@@ 5.4 Ablation Study
Is our ASR evaluation method reliable?
As discussed in Section 5.1, we argue that using “Print exactly Hacked!” as the injected instruction may introduce a misclassification problem. Therefore, we propose an alternative instruction framed as a reasoning task, such as “What is the answer to 1000 + 555?”. However, to ensure reliability, it is crucial that the evaluation dataset contains no sample whose answer contains “1555”. To verify this, we test ASR in the absence of any prompt injection attacks. Table 6 demonstrates the reliability of our evaluation method. For direct prompt injection attacks, the ASR is 0.0, indicating the reliability of our evaluation method. For indirect attacks, only one sample’s answer contains “1555”. Given the significant margin between our results and the baselines, this slight bias can be disregarded. Therefore, we conclude that our evaluation method is reliable.
Can our method defend against the GCG attack?
Beyond prompt-engineering-based attacks, we also examine the effectiveness of our defense methods against the GCG attack. We implement the GCG attack using Llama3, following the implementation by Chen et al. (2024), and construct the suffix with knowledge of the conversation template. Since the GCG suffix is transferable in jailbreak attacks Zou et al. (2023), we also test its transferability in prompt injection attacks using Llama3.1. Table 5 presents the defense results. The first notable observation is that the suffix transferability is limited. Without any defense, the ASR remains relatively low, similar to a naive attack. While our defense methods based on ignore attack, escape attack, and fake completion attack show less satisfactory performance, the defense method based on "Fakecom-t" still effectively mitigates the attack.
@@
Table 5: The performance of the defense methods against the GCG attack. All results are reported in %.
@@ A.1 Defense Baselines
Incontext Wei et al. (2023).
This method provides poisoned data content along with the correct answers, allowing LLMs to learn through in-context learning. An example is shown in Fig 7.
@@ A.3 Dataset Additional Information
Scenario llama3 llama3.1 qwen2
direct 0.0 0.0 0.0
indirect 0.05 0.05 0.05
Table 6: The ASR results when no attack and defense is applied to the input data. All results are reported in %.

### v2

@@ Abstract
With the advancement of technology, large language models (LLMs) have achieved remarkable performance across various natural language processing (NLP) tasks, powering LLM-integrated applications like Microsoft Copilot. However, as LLMs continue to evolve, new vulnerabilities, especially prompt injection attacks arise. These attacks trick LLMs into deviating from the original input instructions and executing the attacker’s instructions injected in data content, such as retrieved results. Recent attack methods leverage LLMs’ instruction-following abilities and their inabilities to distinguish instructions injected in the data content, and achieve a high attack success rate (ASR). When comparing the attack and defense methods, we interestingly find that they share similar design goals, of inducing the model to ignore unwanted instructions and instead to execute wanted instructions. Therefore, we raise an intuitive question: Could these attack techniques be utilized for defensive purposes? In this paper, we invert the intention of prompt injection methods to develop novel defense methods based on previous training-free attack methods, by repeating the attack process but with the original input instruction rather than the injected instruction. Our comprehensive experiments demonstrate that our defense techniques outperform existing defense approaches, achieving state-of-the-art results.
@@ 1 Introduction
We conduct comprehensive experiments to evaluate the effectiveness of our defense methods against various prompt injection attack methods. The results demonstrate that our methods outperform existing training-free defense approaches and are even comparable to fine-tuning-based methods, providing more robust protection. Notably, the defense method based on the most effective attack technique performs the best, reducing the attack success rate (ASR) to nearly zero in certain scenarios. Our contributions are summarized as follows:
@@
  • We develop prompt injection defense methods based on attack strategies, which demonstrate greater effectiveness compared to existing baselines.
@@ 5.1 Experimental Settings
For the security metric, we follow the evaluation protocol of Chen et al. (2024), using the attack success rate (ASR) to assess the effectiveness of the defense methods. We detecte if the answer to the injected instruction appears in the generated response. For the utility metric, we use accuracy to evaluate the potential negative impact of defense methods on model performance. Specifically, we employ the filtered QA dataset Li et al. (2023), which is not attacked and includes the defense mechanism. We request the LLMs to answer the questions and verify whether the correct (golden) answers appear in the responses.
@@ 5.2.2 Defense Baselines
For a fair comparison, we select existing training-free defense methods as baselines. Specifically, we select Sandwich san (2023), Instructional ins (2023), Reminder Yi et al. (2023), Isolation Willison (2023), Spotlight Hines et al. (2024) for comparison. More details about the baselines can be found in Appendix A.1.
@@ 5.3.3 Model Utility
None 78.05 77.10 76.60
Sandwich 80.80 79.50 77.35
Instructional 77.30 79.30 75.35
Reminder 77.20 78.05 76.05
Isolation 78.10 78.25 77.10
Spotlight 76.40 77.90 78.25
Ours-Ignore 78.55 79.60 77.60
Ours-Escape 79.40 79.85 80.40
Ours-Fakecom 80.75 80.45 81.30
Ours-Fakecom-t 79.40 80.45 77.40
Table 3: The model utility when applied with different defense methods. All the results are reported in %.
@@ 5.4 Ablation Study
Can our methods be extended to the closed-source models?
To further validate the effectiveness of our methods, we apply our methods to the closed-source models "GPT-3.5-Turbo" Jiang et al. (2023) and "GPT-4o-Latest" Hurst et al. (2024). Because we cannot change the conversation template, we only compare our methods based on "Ignore attack" and "Fakecom attack" with the baselines against direct prompt injection attack. Table 4 shows the results. From the table we can find out that our methods are also effective on closed-source models, surpassing the previous defense baselines. What’s more, comparing the defense performance of the two models with our defense methods reveals that stronger model is more suitable to our methods, making our methods more applicable.
Defense Methods GPT-3.5-Turbo GPT-4o-Latest
Naive Ignore Escape Fakecom Combined Naive Ignore Escape Fakecom Combined
None 32.69 50.48 32.69 88.46 87.50 65.86 92.78 63.46 100.00 100.00
Sandwich 13.94 17.30 8.65 4.32 42.30 15.86 29.32 10.09 5.76 37.98
Instructional 25.00 34.61 26.92 44.23 71.63 24.51 18.26 28.36 62.01 42.78
Reminder 11.05 10.57 10.96 9.61 26.92 14.90 27.88 17.30 89.42 82.69
Isolation 22.59 39.42 24.51 43.26 77.40 52.40 83.17 52.88 94.71 99.51
Spotlight 16.34 31.73 13.46 15.38 71.15 19.71 45.67 15.38 47.11 68.75
Ours-Ignore 2.88 3.36 1.44 0.48 4.32 0.90 0.90 0.40 0.0 0.0
Ours-Fakecom 5.57 12.01 1.44 0.0 14.90 7.21 2.88 5.76 0.90 7.21
Table 4: The results of our defense methods compared with defense baselines applied on closed-source models. Bold indicates the best performance. All results are reported in %.
Can our methods defend against the gradient-based attack?
Beyond prompt-engineering-based attacks, we also examine the effectiveness of our defense methods against the gradient-based attack. We implement direct prompt injection attack with GCG attack method using Llama3, following the implementation by Chen et al. (2024), and construct the suffix with knowledge of the conversation template. Since the GCG suffix is transferable in jailbreak attacks Zou et al. (2023), we also test its transferability in prompt injection attacks using Llama3.1. Table 6 presents the defense results. The first notable observation is that the suffix transferability is limited. Without any defense, the ASR remains relatively low, similar to a naive attack. Comparing with the baselines, our defense methods more effectively mitigate the attack. The method based on “Fakecom-t” is the most effective, reducing the ASR to below 10%.
@@
Can our methods compare with the fine-tuning-based methods?
In previous introduction, we argue that fine-tuning-based methods require significant computational resources. But the fine-tuning-based methods are more effective than previous prompt-engineering-based methods. Therefore, we compare our methods with StruQ Chen et al. (2024), which incorporates prompt injection attack methods into the clean data for fine-tuning. We incorporate the "Naive attack" and the "Ignore attack" respectively for evaluation. Table 7 shows the results. It’s obvious that the ability of StruQ to generalize to the unknown attacks is not satisfactory. Because "Ignore attack" is a part of "Combined attack", "StruQ-Ignore" can defend against "Combined attack" successfully. The generalization ability of our methods is much better, effectively defending against different attacks.
@@
Table 6: The performance of the defense methods against the gradient-based attack. Bold indicates the best performance. All results are reported in %.
@@ A.3 Dataset Additional Information
Defense Methods Naive Ignore Escape Fakecom Combined
None 10.55 53.35 88.25 75.30 86.00
StruQ-Naive 0.50 0.60 2.20 35.55 27.30
StruQ-Ignore 0.05 0.05 8.00 35.70 0.05
Ours-Ignore 0.05 0.35 0.30 0.10 1.35
Table 7: Defense performance of our methods and the fine-tuning method StruQ. "StruQ-Naive" means StruQ incorporates the "Naive attack" for fine-tuning. Bold indicates the best performance. All results are reported in %.